Week of 070115

Open Actions from last week:

Chair: H.Renshall

Gmod: R.Naujikas

Smod: I.Reguero

Monday:

Log: Nothing

New Actions:

Discussion: FTA_wrong alarms of last week were due to a misconfiguration in old servers which have now been switched off. Only FTS traffic out of CERN continues to be CMS transfers to FNAL at a steady 200 MB/s.

Tuesday:

Log:

New Actions:

Discussion: JvE - castorgrid is being upgraded this week - will be transparent. JC - will take an lfc out of production to instrument it to trace the crashes seen before Xmas. HRR to check with lhcb when we can reinstall volhcb02 (has 2 jbod system disks). HRR to follow up on lxfs6001 pacct file filling up. Multi-VO tests will be restarted this week. Waiting for answers from RAL, CNAF and FZK and CMS. ALICE has already restarted to IN2P3 and FZK.

Wednesday

Log:

New Actions: HRR/JvE to organise refurbishment of lxdpm01. Uli to reinstall volhcb02 Monday morning (if lhcb agree).

Discussion: lxfs6001 (/var full) is now understood (an old twverify bug) and can be used to replace lxdpm01 which needs refurbishing by Elonex. lhcb agree that volhcb02 can be reinstalled (has jbod system disks). The proceure for updating the firmware of the 118 Transtec Western Digital disk servers (4000 disks) is ready. The plan is to perform this in February. It requires the disk servers to be taken out of production so will be difficult to schedule.

Thursday

Log: Network firewall changes completed (7 to 8 am).

New Actions: ITRAC integration testbed to have 2 patches applied tomorrow morning 10 to 11. Tape robot intervention till 14.00 so slow mounting for some tapes.

Discussion: GSSDATLAS now transferring to FZK at 80 MB/s, ALICE to FZK at 100 MB/s and to IN2P3 at 40 MB/s. HRR to discuss FTS shares with GM.

Friday

Log:

New Actions:

Discussion: GSSDATLAS was running at 100 mb/s to fzk but stopped at 03.00 GMT. Many gridbdii high load alarms after reinstall of bdii108 to convert from jbod to raid-1. Reason was reinstall does not automatically run yaim configuration on grid servers- needs to be done by hand. Fixed by 10.00. bdii107 remains to be done.

Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2007-02-02 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback